Problem 1 Consider the sinusoidal signal x(t) shown below. 10 8 4 2 4 6 8 -10 0.2 0.4 0.6 0.8 t (seconds) (i) What is the amplitude of this signal? ii) What is the period of this signal (in seconds)? (iii) What is the frequency of this signal (in Hertz)? (iv) Write an equation for this signal.
Answer #1

a)

A = 8

======

b)

time period of wave,

T = 0.2 seconds

=======

c)

f = 1/ T = 1/ 0.2 = 5 Hz

========

d)

x = A sin ( pi/2 + 2 pi f t)

x = 8 cos ( 2* 3.14* 5* t)

x = 8 cos ( 31.4 t)

========

do comment in case any doubt, will reply for sure.. Goodluck